OpenStudy (waheguru):

Sales Goals Alexis works part-time at a clothing store. She is paid an hourly rate of $10.25/h and also earns a commission of 3.5% of her total weekly sales. Alexis works at the store 12 hours a week. If Alexis’s goal is to earn $150 every week, what do her total weekly sales need to be? Show your work. i did it i just wanna know if in rite

OpenStudy (anonymous):

So, this is the equation I came up with - same as yours? (X equals weekly sales) \[10.25*12+.035*X=150\]

OpenStudy (waheguru):

yep

OpenStudy (anonymous):

Sorry, got busy there - well, it looks like you've got the right answer then. 10.25*12 = 123 So... 123 + .035X = 150 .035X = 150 -123 .035X = 27 X = 771.42 (I'd round it to 772 so she needs to sell $772 to make her $150 total)